OpenStudy (poopsiedoodle):

What effect does the length of the string of a pendulum have on the pendulum's period?

OpenStudy (anonymous):

\[\large T \approx 2\pi \sqrt\frac{L}{g}\] From this equation you can see that as the length of the 'rod' increases than the period will also increase

OpenStudy (poopsiedoodle):